### v1

# Variational Calculation of the Hyperfine Stark Effect in Atomic 87Rb, 133Cs, and 169Tm

Abstract
An electronically variational approach to the calculation of atomic hyperfine structure transition energies under the influence of static external electric fields is presented. The method avoids the calculation of intermediate atomic states entirely and requires only the wavefunctions of the electronic states involved in the respective hyperfine levels. These wavefunctions are obtained through relativistic general-excitation-rank configuration interaction theory. The method also enables for calculations on atoms with the most complicated of shell structures.
The initial applications include 87Rb and 133Cs where very good agreement of the approach with literature results is established. For 169Tm that is used in the development of atomic clocks the differential static electric dipole polarizability between ground levels $J=\frac{7}{2}$ and $J=\frac{5}{2}$ is calculated to be $\Delta\alpha=-0.23\pm 0.11$ a.u. The hyperfine Stark coefficient for the hyperfine levels belonging to the ground term with $J=\frac{7}{2}$ is found to be $k=(1.3\pm 1.0)\times 10^{-13}$ [Hz/((V/m)2)]. This coefficient is several orders of magnitude smaller than the corresponding coefficients in 87Rb and 133Cs.
I Introduction
Among the various effects ludlow_rmp_2015 contributing to the frequency uncertainty of an atomic clock the blackbody radiation (BBR) shift plays a major role Itano_Wineland_Blackbody1982 ; porsev_derevianko_PRA2006 . The BBR shift, a systematic environmental perturbation, can be approximately related to the coefficient of the static electric-field shift of the clock transition Angstmann_Dzuba_Flambaum_PRL2006 ; Kozlov_Dzuba_Flambaum_PRA2014 . The latter coefficient, in turn, describes the difference between the static electric dipole polarizabilities of the respective clock-transition states Safronova_BBR_2009 . Their calculation typically requires a summation over a complete (bound and scattering) set of intermediate atomic states which can be problematic and has in the past led to disagreements (Angstmann_Dzuba_Flambaum_PRA2006 and references therein).
If the clock transition involves atomic hyperfine states then the static electric-field shift of those hyperfine levels in the external electric field schwartz_SPT_1959 ; Fortson_Ramsey_1964 needs to be determined. The present approach takes an external electric field into account variationally – i.e., to infinite order expressed in the language of perturbation theory – in the optimization of the atomic wavefunction. In this way, the summation over intermediate states Angstmann_Dzuba_Flambaum_PRA2006 ; beloy_derevianko_PRL2006 ; Safronova_UI_MS_2009 is avoided and only the field-dependent target-state electronic wavefunction of the atom has to be determined. The static Stark shift of a hyperfine clock transition then results from a simple expectation value of this wavefunction over the magnetic hyperfine operator.
Difficulties in addressing atomic quantum states with several open electronic shells (i.e., containing several unpaired electrons) and/or atomic states with a few holes in otherwise filled shells have been reported Safronova_BBR_2009 . The present approach and method has no such particular difficulties and the wavefunctions for more complicated quantum states of atoms can be obtained straightforwardly. This has recently been demonstrated in the calculation of electronic electric quadrupole moments of the thulium atom Fleig_PhysRevA.107.032816 which comprises an open $f$ shell in its electronic ground state.
Indeed, hyperfine-level transitions in the thulium atom have recently been used as clock transitions in the design of an optical atomic clock with unusually low sensitivity to the BBR NatComm_Tm_2019 ; PRA_Tm_2016 ; PRA_Tm_2020 . The relevant BBR frequency shift has been experimentally estimated NatComm_Tm_2019 to be a few orders of magnitude smaller than the corresponding shift in clock transitions of other neutral atoms. The uncertainty of this measurement has been given as around $50$% AIP_Tm_2020 for the electric component of the BBR. An extensive theoretical study of the thulium atom with relevance to its use as an atomic clock has been presented very recently fritzsche_Tm_2024 , but the BBR or hyperfine Stark shifts have not been addressed in that work.
The paper is organized as follows. In section II the theoretical approach is laid out in detail and a comparison with the common perturbative approaches to the calculation of hyperfine Stark shifts is drawn. In the following section III the hyperfine Stark coefficients for three atoms of relevance to atomic-clock research are presented using the present method. In addition, the differential static dipole polarizability relevant to the clock transition in 169Tm is calculated using a finite-field approach that has earlier been applied to the calculation of spin-orbit resolved electric dipole polarizabilities in atomic states fleig_halo ; fleig_pol13 . In the final section IV conclusions from the presented work are drawn.
II Theory
II.1 Static Hyperfine Stark Shift
We suppose a hyperfine doublet of states denoted with total angular-momentum quantum numbers $F_{u}$ (upper) and $F_{l}$ (lower) is subjected to a static external electric field $E_{\text{ext}}$. Treating the hyperfine interaction to first order in perturbation theory, the associated field-dependent transition energy can be written as
$\Delta\varepsilon(E_{\text{ext}})=\varepsilon_{J_{u},{M_{J}}_{u}}(E_{\text{ext}})+\varepsilon_{F_{u}}(E_{\text{ext}})-\big{(}\varepsilon_{J_{l},{M_{J}}_{l}}(E_{\text{ext}})+\varepsilon_{F_{l}}(E_{\text{ext}})\big{)}.$ (1)
where $J$ denotes total electronic angular momentum associated with the respective hyperfine level. $\varepsilon_{F}$ is thus the hyperfine energy relative to the respective electronic reference energy $\varepsilon_{J,M_{J}}$. Note that generally $J_{u}\neq J_{l}$.
This transition energy can be related to a transition frequency which is also field dependent:
$\nu(E_{\text{ext}})=\frac{\Delta\varepsilon(E_{\text{ext}})}{h}$ (2)
where $h$ is Planck’s constant. The shift of this transition frequency due to $E_{\text{ext}}$ is in the present evaluated by using a finite but small $E$ field and a zero-field calculation. Then
$\delta\nu(E_{\text{ext}})=\nu(E_{\text{ext}})-\nu(E_{\text{ext}}=0).$ (3)
Following the definition in Ref. Safronova_BBR_2009 with according modifications and taking the electronic degrees of freedom into account the Stark coefficient $k$ as a function of this frequency shift is expressed as
$k=\frac{\delta\nu(E_{\text{ext}})}{E_{\text{ext}}^{2}}$ $=$ $\frac{1}{hE_{\text{ext}}^{2}}\left[\big{(}\varepsilon_{J_{u},{M_{J}}_{u}}(E_{\text{ext}})+\varepsilon_{F_{u}}(E_{\text{ext}})-\varepsilon_{J_{l},{M_{J}}_{l}}(E_{\text{ext}})-\varepsilon_{F_{l}}(E_{\text{ext}})\big{)}\right.$ (5)
$\hskip 42.67912pt\left.-\big{(}\varepsilon_{J_{u},{M_{J}}_{u}}(0)+\varepsilon_{F_{u}}(0)-\varepsilon_{J_{l},{M_{J}}_{l}}(0)-\varepsilon_{F_{l}}(0)\big{)}\right]$
$=$ $\frac{1}{hE_{\text{ext}}^{2}}\left[\big{(}\varepsilon_{J_{u},{M_{J}}_{u}}(E_{\text{ext}})-\varepsilon_{J_{l},{M_{J}}_{l}}(E_{\text{ext}})\big{)}+\big{(}\varepsilon_{F_{u}}(E_{\text{ext}})-\varepsilon_{F_{l}}(E_{\text{ext}})\big{)}\right.$
$\hskip 42.67912pt\left.-\big{(}\varepsilon_{J_{u},{M_{J}}_{u}}(0)-\varepsilon_{J_{l},{M_{J}}_{l}}(0)\big{)}-\big{(}\varepsilon_{F_{u}}(0)-\varepsilon_{F_{l}}(0)\big{)}\right]$
and can, therefore, be calculated if the field-dependent level energies are known.
At this point a distinction between two cases can be made:
  1. 1.
    $J_{u}\neq J_{l}$, i.e., the two respective hyperfine levels belong to different electronic states. Then the Stark coefficient is stongly dominated by the electric polarizability difference between the different electronic wavefunctions and it can be written as
    $k\approx\frac{1}{hE_{\text{ext}}^{2}}\left[\big{(}\varepsilon_{J_{u},{M_{J}}_{u}}(E_{\text{ext}})-\varepsilon_{J_{l},{M_{J}}_{l}}(E_{\text{ext}})\big{)}-\big{(}\varepsilon_{J_{u},{M_{J}}_{u}}(0)-\varepsilon_{J_{l},{M_{J}}_{l}}(0)\big{)}\right]$ (6)
  2. 2.
    $J_{u}=J_{l}$, i.e., the two respective hyperfine levels belong to the same electronic states. In this case all of the electronic energies in the expression (5) cancel pairwise and the Stark coefficient is
    $k=\frac{1}{hE_{\text{ext}}^{2}}\left[\big{(}\varepsilon_{F_{u}}(E_{\text{ext}})-\varepsilon_{F_{l}}(E_{\text{ext}})\big{)}-\big{(}\varepsilon_{F_{u}}(0)-\varepsilon_{F_{l}}(0)\big{)}\right]$ (7)
II.2 Field-Dependent Hyperfine Level Energy
We will now be concerned with the above second case. A field-dependent hyperfine energy is related to the magnetic hyperfine constant $A$ as (see bethe_hypfin , p. 110)
$\varepsilon_{F}(E_{\text{ext}})\approx\frac{1}{2}\,\left[F(F+1)-I(I+1)-J(J+1)\right]\,A(E_{\text{ext}})$ (8)
where the quantum numbers $I$ and $J$ refer to nuclear spin and to total electronic angular momentum, respectively, of the state in question. The relationship Eq. (8) is not exact because the presence of the external field lifts the full rotational symmetry of the atom and thus also the strict validity of total angular momentum as a good quantum number. However, since the applied fields are very small (see below) the relation is still approximately correct.
Thus, if the field-dependent hyperfine constant $A(E_{\text{ext}})$ can be determined then also the Stark coefficient $k$. Following the implementation in Refs. Fleig2014 ; Fleig_Skripnikov2020 $A(E_{\text{ext}})$ is calculated as an expectation value
$A(E_{\text{ext}})=\left<\hat{H}_{\text{HF}}\right>_{\psi(E_{\text{ext}})}$ (9)
with the one-body hyperfine Hamiltonian
$\hat{H}_{\text{HF}}=-\frac{\mu[\mu_{N}]}{2cIm_{p}M_{J}}\,\sum\limits_{i=1}^{n}\,\left(\frac{\boldsymbol{\alpha}_{i}\times{\bf{r}}_{i}}{r_{i}^{3}}\right)_{z}$ (10)
over a field-dependent electronic wavefunction $\psi(E_{\text{ext}})$ where $n$ is the number of electrons, $\boldsymbol{\alpha}$ is a Dirac matrix, $\mu$ is the nuclear magnetic moment [in nuclear magnetons], $\frac{1}{2cm_{p}}$ is the nuclear magneton in a.u. , $m_{p}$ is the proton rest mass, and ${\bf{r}}$ is the electron position operator.
$\psi(E_{\text{ext}})$ is obtained by solving
$\hat{H}({{E}_{\text{ext}}})\big{|}\psi(E_{\text{ext}})\big{>}=\varepsilon({{E}_{\text{ext}}})\big{|}\psi(E_{\text{ext}})\big{>}$ (11)
with $\varepsilon({{E}_{\text{ext}}})$ the field-dependent energy eigenvalue and $\hat{H}({{E}_{\text{ext}}})$ the Dirac-Coulomb (DC) Hamiltonian including the interaction term with the external field:
$\hat{H}({{E}_{\text{ext}}})$ ::=:= $\hat{H}^{\text{Dirac-Coulomb}}+\hat{H}^{\text{Int-Dipole}}$ (12)
$=$ $\sum\limits^{n}_{j}\,\left[c\,\boldsymbol{\alpha}_{j}\cdot{\bf{p}}_{j}+\beta_{j}c^{2}-\frac{Z}{r_{jK}}{1\!\!1}_{4}\right]+\sum\limits^{n}_{k>j}\,\frac{1}{r_{jk}}{1\!\!1}_{4}+\sum\limits_{j}\,{\bf{r}}_{j}\cdot{\bf{E}_{\text{ext}}}\,{1\!\!1}_{4}$
${\bf{E}_{\text{ext}}}=E_{z}{\bf{e}}_{z}$ is uniform in space, the indices $j,k$ run over $n$ electrons, $Z$ is the proton number with the nucleus $K$ placed at the origin, and $\boldsymbol{\alpha},\beta$ are standard Dirac matrices. $E_{z}$ is not treated as a perturbation but included a priori in the variational optimization DIRAC_JCP of the wavefunction, $\psi(E_{z})$.
Technically, $\psi(E_{z})$ is a configuration interaction (CI) vector knecht_luciparII built from Slater determinants over field-dependent 4-spinors. This vector is represented as
[…]
where ${\cal{F}}^{t}(M,n)$ is the symmetry-restricted sector of Fock space with $n$ electrons in $M$ four-spinors, ${\cal{S}}=a^{\dagger}_{i}a^{\dagger}_{j}a^{\dagger}_{k}\ldots$ is a string of spinor creation operators, ${\overline{\cal{T}}}=a^{\dagger}_{\overline{l}}a^{\dagger}_{\overline{m}}a^{\dagger}_{\overline{n}}\ldots$ is a string of creation operators of time-reversal transformed spinors. The determinant expansion coefficients $c_{(\alpha,J,M_{J}),K}$ are generally obtained as described in refs. fleig_gasci ; fleig_gasci2 by diagonalizing the Dirac-Coulomb Hamiltonian.
II.3 Comparison with Perturbation Theory
In order to compare the present method with the perturbative approaches typically used in the literature, it is assumed that a zeroth-order CI problem for the target state $\psi_{0}$ is solved where $\hat{V}=\hat{H}^{\text{Int-Dipole}}=\sum\limits_{j}\,{\bf{r}}_{j}\cdot{\bf{E}_{\text{ext}}}\,{1\!\!1}_{4}$ is not included in the Hamiltonian $\hat{H}^{(0)}$:
$\hat{H}^{(0)}\big{|}\psi_{0}^{(0)}\big{>}=\varepsilon_{0}\big{|}\psi_{0}^{(0)}\big{>}$ (14)
Expanding the wavefunction $\left|\psi_{0}\right>$ for the target state into a perturbation series where $\hat{V}$ is the perturbation and retaining only non-zero terms yields
$\left|\psi_{0}\right>$ $=$ $\left|\psi_{0}^{(0)}\right>+\sum\limits_{k\neq 0}\,\frac{V_{k0}}{\varepsilon_{0}-\varepsilon_{k}}\left|\psi_{k}^{(0)}\right>-\frac{1}{2}\sum\limits_{k\neq 0}\,\frac{|V_{k0}|^{2}}{(\varepsilon_{0}-\varepsilon_{k})^{2}}\left|\psi_{0}^{(0)}\right>$ (15)
$+\sum\limits_{k,l\neq 0}\,\frac{V_{kl}V_{l0}}{(\varepsilon_{0}-\varepsilon_{k})(\varepsilon_{0}-\varepsilon_{l})}\left|\psi_{k}^{(0)}\right>+{\cal{O}}(\hat{V}^{3})$
where $V_{kl}=\left<\psi_{k}^{(0)}|\hat{V}|\psi_{l}^{(0)}\right>$ and $\varepsilon_{m}$ is an unperturbed energy. If terms ${\cal{O}}(\hat{V}^{3})$ are omitted from Eq. (15) and the resulting truncated expansion is used to evaluate the expectation value $\psi_{0}\right|\hat{H}_{\text{HF}}\left|\psi_{0}\right>$ then the third-order expressions from Refs. schwartz_SPT_1959 and Angstmann_Dzuba_Flambaum_PRA2006 arise. In the present work, however, ${\cal{O}}(\hat{V}^{3})$ is implicitly included in Eq. (11) which leads to the presence of all orders of the external electric field in Eq. (9) and, therefore, also in Eq. (5) defining the hyperfine Stark coefficient.
III Applications and results
III.1 Technical Details
III.1.1 External Electric Field
The external electric field $E_{z}$ has to be chosen small enough to assure that $k$ as defined in Eq. (5) does not vary with $E_{z}$, but also large enough to ensure that the field-dependent change in $\psi(E_{z})$ is sufficiently greater than the convergence threshold chosen for the wavefunction. Typically, this is achieved for $E_{z}=10^{-3}$ a.u.  which is the value used in the present applications. In addition, a region around $E_{z}=10^{-3}$ a.u.  has been explored which allows for an estimate of the uncertainty in $k$ due to deviations from quadratic dependency of the field-dependent total energy.
III.1.2 Atomic Basis Sets
For all three atoms Gaussian basis sets are used. The Rb atom is described by the uncontracted aug-cc-pwCVQZ-X2C set from the EMSL library EMSL-basis2019 that has been singly densified in the $s$ and $p$ spaces and augmented by one dense and one diffuse function as described in Ref. Hubert_Fleig_2022 . The full uncontracted set then comprises $77s,59p,21d,5f,3g$ functions. The Cs atom is described by the uncontracted Dyall vQZ basis dyall_s-basis with $5s5p6s6p$ correlating exponents added. Like for Rb the $s$ and $p$ spaces have been densified once and augmented. The final uncontracted set comprises $75s,61p,19d,3f,1g$ functions.
For the thulium atom three different basis sets dyall_4f are employed: 1) The uncontracted Dyall vTZ set including all $[4f/6s/5s,p,d]$ correlating and $4f$ dipole-polarizing functions adding up to $[30s,24p,18d,13f,4g,2h]$ functions (in the following denoted as TZ). 2) The uncontracted Dyall cvQZ basis including all $[4f/6s/5d]$ correlating and $4f$ dipole-polarizing functions comprised by a total of $[35s,30p,19d,16f,6g,4h,2i]$ functions (QZ) and 3) The latter basis set with the $s$ and $p$ spaces simply densified and augmented as described in Ref. Hubert_Fleig_2022 , yielding $[71s,61p,19d,16f,6g,4h,2i]$ primitive functions (QZ+). The densification and augmentation in the $s$ and $p$ spaces serves to produce a highly accurate basis set for describing the $s$-$p$ mixing which is predominantly responsible for the energy shifts when an external electric field is applied to hyperfine states.
III.1.3 Atomic Wavefunctions
All atoms considered here have an odd number of electrons. In addition, the interaction with $E_{z}$ breaks the full rotational symmetry of the atom which means that only $M_{J}$ remains as an exact electronic quantum number (for $\psi(E_{\text{ext}})$ in Eq. (11)). Atomic wavefunctions are calculated in the $\left|M_{J}\right|=\Omega=\frac{1}{2}$ irreducible representation of the $C_{\infty v}^{*}$ double point group. The target atom is placed at the origin of the reference frame and a ghost atom with neither electronic basis functions nor nuclear charge is placed at a finite distance along the axis of the external $E$ field in order to allow for the inclusion of the external-field Hamiltonian.
For solving Eq. (11) the KRCI module knecht_luciparII of the DIRAC program package DIRAC_JCP is used. In a first step the Dirac-Coulomb-Hartree-Fock equations are solved where the Hamiltonian in Eq. (12) is employed. This model will be abbreviated as DCHF. The atomic spinors are optimized by diagonalizing a Fock operator where a fractional occupation of $f=\frac{m}{n}$ per spinor in the defined valence shells is used. Here $m$ is the number of electrons and $n$ is the number of spinors the respective shell comprises. These spinors are thus obtained for the electric potential of the neutral atom.
Acronyms are used for brevity in defining atomic correlated wavefunctions from the second step. As an example, SDT9_10au stands for Single, Double, Triple replacements relative to the DCHF reference state where 9 electrons occupying the outermost shells in the DCHF reference state are taken into account in the correlation expansion and the complementary space of virtual spinors is truncated at 10 a.u. An acronym SD8_SDT9 means that up to two holes are allowed in the shells occupied by 8 electrons and up to triple excitations into the virtual spinors are allowed from the combined shells occupied by the 9 electrons in the reference state.
III.1.4 Finite-Field (FF) Method
Static dipole polarizabilities $\alpha_{J}$ are calculated by fitting finite-field electronic energies for four field points with $E\in\{0,0.00025,0.0005,0.001\}$ a.u. to a polynomial and extracting the second derivative of the fitted function at zero field which is proportional to the static dipole polarizability, see also Refs. fleig_halo ; fleig_pol13 .
III.2 87Rb
For the electronic ground state corresponding to the valence configuration $4s^{1}$ the total electronic angular momentum quantum number is ${J=1/2}$. The considered isotope has $I=3/2$ and the resulting hyperfine quantum levels are denoted as $F_{u}=2$ and $F_{l}=1$. The fractional occupation in the DCHF calculation is $f=1/2$.
Results for the hyperfine Stark coefficient are shown in Table 1 for various electronic-structure models and are compared with experimental and theoretical literature results. The result in mean-field approximation (DCHF) differs from the experimental result by about $20$% and lowest-order electron correlation effects from the $3s,3p,4s$ Rb shells even increase this discrepancy (model SD9). Including virtual spinors up to $30$ a.u.  does not lead to a significant change, so for the computationally most expensive models the virtual cutoff is set to $10$ a.u.
Upon including combined triple excitations from the $3s,3p$ shells and the $4s$ valence shell (model SD8_SDT9) a strong correction to $k$ is obtained that, however, seems to overshoot the exact value. The inclusion of full triple excitations (model SDT9), i.e., including those from the $3s,3p$ shells, proves to be crucial in obtaining an accurate value for $k$. Excellent agreement with the result obtained by Safronova et al. Safronova_BBR_2009 is achieved at this level of calculation.
Table 1: Stark hyperfine coefficient $k$ for 87Rb with nuclear magnetic moment $\mu=2.75131\mu_{N}$ ISOLDE_BWeffect_1993 .
Model $k$ [$10^{-10}$ Hz/((V/m)2)]
DCHF $-1.546$
SD9/10au $-1.662$
SD9/30au $-1.669$
SD8_SDT9/30au $-1.202$
SDT9/10au $-1.265$
SDT9/30au $-1.271$
Exp. Mowat_Li_1972 $-1.23(3)$
Safronova et al. Safronova_BBR_2009 $-1.272$
As a general effect, the hyperfine constant of a given atom in an electronic $S$ state diminishes when an external $E$ field is included. This observation is explained by the fact that the $E$ field partially shifts spin density from $s$ wave to $p$ wave character in the atomic ground state (through $s$-$p$ mixing in the polarized atom), thus reducing the hyperfine interaction. On the other hand, electron correlation effects increase the hyperfine constants of $S$ state atoms, which means that effectively spin density residing in $s$ spinors is increased by allowing for the mixing in of determinants with open-shell character.
III.3 133Cs
This important isotope of cesium serves as a second test case. For the electronic ground state corresponding to the valence configuration $5s^{1}$ the total electronic angular momentum quantum number is ${J=1/2}$. The considered isotope has $I=7/2$ and the resulting hyperfine quantum levels are denoted as $F_{u}=4$ and $F_{l}=3$. The fractional occupation in the DCHF calculation is $f=1/2$. Results for 133Cs are compiled in Table 2.
As in the case of 87Rb the mean-field result for $k$ is too large on the absolute, compared with the cited reference values, by over $40$%. Again, the inclusion of lowest-order electron correlation effects even increases this deviation. It is here in addition and for the model SD9 shown that the consideration of virtual spinors of very high energy (up to $1000$ a.u. ) in the wave-function expansion does not affect the results significantly. Also as for 87Rb combined triple excitations from the outermost shells yield the decisive correction to $k$. Including full triple excitations (model SDT9/30au) yields a result that deviates from the most accurate literature results by roughly $+9$%. Accounting for combined quadruple excitations quenches the deviation to $-7$% where again, as in the Rb atom, the correction slightly overshoots the exact value. This latter model comprises about $1.9$ billion ($10^{9}$) terms in the CI expansion.
Table 2: Stark hyperfine frequency shift $k$ for 133Cs ($5s_{1/2}$), $I=3.5$, $E_{z}=0.001$ a.u., hyperfine level quantum numbers are $F_{u}=4$, $F_{l}=3$; basis set Dyall-cvQZ+sp
Model $k$ [$10^{-10}$ Hz/((V/m)2)]
DCHF $-3.220$
SD9/10au $-3.375$
SD9/30au $-3.379$
SD9/1000au $-3.384$
SD8_SDT9/10au $-2.308$
SD8_SDT9/30au $-2.311$
SDT9/10au $-2.468$
SDT9/30au $-2.469$
SDT8_SDTQ9/10au $-2.112$
Exp.Simon_PRA1998 $-2.271(4)$
Exp.Godone_PRA2005 $-2.05(5)$
Safronova et al.beloy_derevianko_PRL2006 $-2.271(8)$
Angstmann et al.Angstmann_Dzuba_Flambaum_PRA2006 $-2.26(2)$
In view of the observed trends and their magnitude it is expected that the remaining small deviation from the most accurate experimental value for the Stark shift can further largely be removed by including full quadruple excitations in the wavefunction expansion. Such a calculation is not impossible but would be very time (and resource) consuming and beside the point. It has here been demonstrated that the presented method yields accurate hyperfine Stark shifts that systematically converge towards reliable reference results.
III.4 169Tm
With confidence in the method established in the aforegoing sections it is now applied to an atom where high-level reference results are not available and which has a more complex electronic structure. In this case the fractional occupation in the DCHF calculation is $f=15/16$ where $m=15$ represents the thirteen $4f$ electrons plus the two $6s$ electrons. This averaging was required to assure proper convergence of the DCHF wavefunction.
III.4.1 ${}^{2}F_{7/2}(F=4)$${}^{2}F_{5/2}(F=3)$ clock transition
The clock transition discussed in Refs. NatComm_Tm_2019 ; PRA_Tm_2016 ; PRA_Tm_2020 comprises a hyperfine component ($F_{l}=4$) of the ground electronic state ${}^{2}F_{7/2}$ and a hyperfine component ($F_{u}=3$) of the first electronically excited state ${}^{2}F_{5/2}$. In this case the Stark coefficient is dominated by the polarizability difference between the two respective electronic states and Eq. (6) applies.
Denoting the static electric dipole polarizability as $\alpha_{J,M_{J}}$ the differential polarizability for the present levels is defined as $\Delta\alpha_{M_{J}}=\alpha_{\frac{5}{2},M_{J}}-\alpha_{\frac{7}{2},M_{J}}$. The relationship with the Stark coefficient can then be written as $-2k_{M_{J}}=\Delta\alpha_{M_{J}}$. The results in Table 3 are calculated using Eq. (6) and $\Delta\alpha_{M_{J}}$ is then derived from those values. In addition, individual static polarizabilities $\alpha_{J}$ are calculated by using the FF method. Generally, the single-field-point results match those from the more accurate FF method up to three significant digits when $E=0.0005$ a.u. is used. With $E=0.001$ a.u. correspondence is up to two digits.
Table 3: Stark frequency shift $k$ calculated according to Eq. (6), ground-state electric polarizability and differential polarizability for 169Tm ($4f_{7/2,5/2}$), $E_{z}=0.001$ a.u., $M_{J}=1/2$
Basis set/Model $k$ [$10^{-9}$ Hz/((V/m)2)] $\Delta\alpha$ [a.u.] $\alpha_{\frac{7}{2}}$ [a.u.]
QZ/DCHF $1.42$ $-0.115$ $186.6$
QZ+/DCHF $1.44$ $-0.115$
QZ/SD15_10au $0.52$ $-0.042$ $161.2$
QZ/SDT15_10au $1.77$ $-0.142$
QZ/SD23_10au $1.34$ $-0.108$ $151.2$
QZ/SD33_10au $1.57$ $-0.126$ $151.8$
QZ/SD41_10au $1.58$ $-0.127$
final $2.83$ $-0.23\pm 0.11$
Experiment $-0.063(30)$ NatComm_Tm_2019 $130\pm 16$ Ma_Tl_pol
Recommended $144\pm 15$ schwerdt_nagle_pol_atoms_2018
The final value for $\Delta\alpha$ is calculated by using the individual result from model QZ/SD41_10au with the largest number of correlated electrons as a base value and adding a correction due to taking into account full triple excitations among the $15$ outermost electrons, formally,
[…]
Interestingly, the present final result for $\Delta\alpha$ differs from the one measured and calculated in Ref. NatComm_Tm_2019 , even if assigned uncertainties are taken into account. Since the differential polarizability amounts to only about $0.2$% of an individual level polarizability, the former is by its nature very difficult to calculate accurately. The cited result for $\Delta\alpha$ from Ref. NatComm_Tm_2019 is also given with a relatively large uncertainty, in part due to uncertainties in employed transition probabilities that have been estimated ibidem.
Nevertheless, the present series of calculations allows for some conclusions on this problem. The major correction to the base value comes from higher excitation ranks in the wavefunction expansion (triples terms) and is negative. The importance of this correction is not surprising, as these higher excitations can be interpreted as the mixing in of excited states fulfilling $\Delta J=0$ into the atomic wavefunctions with defined $J$, thus leading to a differential effect for the $\alpha_{J}$ results. It is, therefore, to be expected that including higher excitation ranks (quadruples etc.) will lead to a more negative value for $\Delta\alpha$ and thus to an even greater discrepancy with the result by Golovizin et al.
There is good correspondence of the present most accurate result with high-level literature results for $\alpha_{J}$, but this agreement does not yield a constraint on the uncertainty of the much smaller differential polarizability $\Delta\alpha$.
III.4.2 ${}^{2}F_{7/2}(F=4)$${}^{2}F_{7/2}(F=3)$ transition
When the hyperfine transition among the two microstates belonging to the same ${}^{2}F_{7/2}$ ground level is considered, Eq. (7) is relevant. Results for the ground-term hyperfine Stark coefficient are compiled in Table 4.
Table 4: Stark hyperfine frequency shift $k$ calculated according to Eq. (5) for 169Tm ($4f_{7/2}$), $I=0.5$, $\mu=-0.2310(15)\mu_{N},E_{z}=0.001$ a.u., hyperfine level quantum numbers for the ground-state transition are $(J=7/2)F_{l}=4$, $(J=7/2)F_{u}=3$. All results are given for the electronic microstates with $M_{J}=1/2$ except where stated otherwise.
Basis set/Model $k$ [$10^{-14}$ Hz/((V/m)2)]
TZ/DCHF $-1.38$
TZ/SD15_6au $-6.36$
TZ/SD15_10au $-6.85$
TZ/SD15_30au $-7.80$
TZ/SD15_60au $-7.69$
QZ/DCHF($M_{J}=1/2$) $-1.41$
QZ/DCHF($M_{J}=3/2$) $-1.20$
QZ/DCHF($M_{J}=5/2$) $-0.57$
QZ/SD15_6au $-5.22$
QZ/SD15_10au $-7.32$
QZ/SD15_30au $-7.76$
QZ/SD15_60au $-7.67$
QZ/SDT15_10au $2.08$
QZ/S8_SD23_10au $4.93$
QZ/SD23_10au $4.98$
QZ/SD23_30au $4.28$
QZ/SD33_10au $4.47$
QZ/SD41_10au $3.94$
QZ+/DCHF $-1.41$
QZ+/SD15_5au $-5.12$
QZ+/SD15_10au $-7.00$
k(final) ${\boldsymbol{13.0\pm 10.0}}$
Concerning atomic basis sets, DCHF results are seen to be converged already for the QZ set. When correlation effects among the $15$ outermost electrons are included the largest densified basis set QZ+ gives a correction of $4-5$% to the hyperfine Stark coefficient which will be considered in the final evaluation. Additional correlation effects, on the other hand, are investigated using the computationally less expensive QZ set.
DCHF results for the $M_{J}=1/2$ microstates show the largest hyperfine Stark effect, whereas the effect of the electric field reduces as $M_{J}$ is increased (QZ/DCHF models). This can be explained as follows. The respective (leading, if a model including electron correlation effects is considered) electronic Slater determinants can be written in terms of occupied spinors in the $4f$ shell as
$\left|J=\frac{7}{2},M_{J}=\frac{1}{2}\right>$ $=$ $\left|4f_{5/2,5/2\ldots-5/2};4f_{7/2,7/2};4f_{7/2,-7/2};4f_{7/2,5/2};4f_{7/2,-5/2};4f_{7/2,3/2};4f_{7/2,-3/2};4f_{7/2,1/2}\right>$
$\left|J=\frac{7}{2},M_{J}=\frac{3}{2}\right>$ $=$ $\left|4f_{5/2,5/2\ldots-5/2};4f_{7/2,7/2};4f_{7/2,-7/2};4f_{7/2,5/2};4f_{7/2,-5/2};4f_{7/2,3/2};4f_{7/2,1/2};4f_{7/2,-1/2}\right>$
$\left|J=\frac{7}{2},M_{J}=\frac{5}{2}\right>$ $=$ $\left|4f_{5/2,5/2\ldots-5/2};4f_{7/2,7/2};4f_{7/2,-7/2};4f_{7/2,5/2};4f_{7/2,3/2};4f_{7/2,-3/2};4f_{7/2,1/2};4f_{7/2,-1/2}\right>.$
[…]
$A_{J,M_{J}}$ $=$ $4f_{j,m_{j}}\right|\hat{H}_{\text{HF}}\left|4f_{j,m_{j}}\right>$ (16)
${\tilde{A}}_{J,M_{J}}$ $=$ ${\tilde{4f}}_{j,m_{j}}\right|\hat{H}_{\text{HF}}\left|{\tilde{4f}}_{j,m_{j}}\right>$ (17)
where $\tilde{X}$ denotes a quantity perturbed by the external electric field. According to Eq. (8) the energy of the lower hyperfine level is
$\varepsilon_{F,J,M_{J}}=\varepsilon_{4,\frac{7}{2},M_{J}}\approx\frac{1}{2}\,\left[20-\frac{3}{4}-\frac{63}{4}\right]\,A_{\frac{7}{2},M_{J}}=\frac{7}{4}A_{\frac{7}{2},M_{J}}$ (18)
and the energy of the upper hyperfine level is
$\varepsilon_{3,\frac{7}{2},M_{J}}\approx\frac{1}{2}\,\left[12-\frac{3}{4}-\frac{63}{4}\right]\,A_{\frac{7}{2},M_{J}}=-\frac{9}{2}A_{\frac{7}{2},M_{J}}$ (19)
Correspondingly, when $E_{\text{ext}}$ is included,
$\tilde{\varepsilon}_{4,\frac{7}{2},M_{J}}\approx\frac{7}{4}\tilde{A}_{\frac{7}{2},M_{J}}$ (20)
$\tilde{\varepsilon}_{3,\frac{7}{2},M_{J}}\approx-\frac{9}{2}\tilde{A}_{\frac{7}{2},M_{J}}$ (21)
Note that $\varepsilon_{4,\frac{7}{2},M_{J}}<\varepsilon_{3,\frac{7}{2},M_{J}}$ since $\tilde{A}_{\frac{7}{2},M_{J}}<0$ ${}~{}\forall M_{J}$ as numerical results show which makes the level with $F=4$ the lower level.
Then the Stark coefficient for states with defined $M_{J}$ results from Eq. (5) as
$k_{M_{J}}$ $=$ $\frac{\tilde{\varepsilon}_{3,\frac{7}{2},M_{J}}-\varepsilon_{3,\frac{7}{2},M_{J}}-\tilde{\varepsilon}_{4,\frac{7}{2},M_{J}}+\varepsilon_{4,\frac{7}{2},M_{J}}}{hE_{\text{ext}}^{2}}$ (22)
$=$ $\frac{1}{hE_{\text{ext}}^{2}}\left[-\frac{9}{2}\left(\tilde{A}_{\frac{7}{2},M_{J}}-A_{\frac{7}{2},M_{J}}\right)-\frac{7}{4}\left(\tilde{A}_{\frac{7}{2},M_{J}}-A_{\frac{7}{2},M_{J}}\right)\right]$
$=$ $\frac{1}{hE_{\text{ext}}^{2}}\left(-\frac{25}{4}\right)\left(\tilde{A}_{\frac{7}{2},M_{J}}-A_{\frac{7}{2},M_{J}}\right)$
Numerical evaluation of this expression shows that indeed $\left|k_{\frac{1}{2}}\right|>\left|k_{\frac{3}{2}}\right|>\left|k_{\frac{5}{2}}\right|$ which in essence is due to the magnitude of $\tilde{A}_{J,M_{J}}-A_{J,M_{J}}$ decreasing with increasing value of $M_{J}$. This, in turn, is readily explained by the fact that the orbital amplitude of the $4f$ wavefunction in field direction ($z$ axis) is maximal for minimal $M_{J}$, thus leading to the greatest field-dependent shift when $M_{J}=\frac{1}{2}$.
Including full triple excitations among the $15$ valence electrons (model SDT15) yields a relatively sizeable positive correction that brings the hyperfine Stark coefficient back close to the DCHF value (however, with opposite sign). It is remarkable that this correction counteracts the correlation effect from single and double excitations. Since the model SDT15 comprises nearly one billion expansion terms it is computationally impossible with the present code to investigate the effects of even higher excitation ranks than full triple replacements.
Accounting for electron correlation effects among and with the eight electrons in the Tm $5s,5p$ shells (model SD23_30au) yields a sizeable positive correction of $+155$% relative to the model SD15 that in addition changes the sign of $k$. It is not surprising that such physical effects are important since the hyperfine structure in an external $E$ field is dominated by the atomic structure in $s$ and $p$ states. In accord with this reasoning, the inclusion of correlation effects with and among the $4d$ electrons (model SD33) results in an only relatively minor correction of about $-10$%. Adding the $4s$ and $4p$ electrons (model SD41)yields a slightly more important contribution of about $-12$%, again in line with the above arguments.
The final value for the thulium hyperfine Stark shift can now be obtained from the above series of calculations. The result with $23$ electrons considered in the correlation treatment and the highest cutoff value for the virtual spinors serves as the base value. This base value is corrected for by the effects of densifying the basis set, correlating inner-shell electrons and taking triple excitations from the $15$ outermost electrons into account. Formally,
[…]
which is given in Table 4. This final result is greater than any of the individual results, both as a plain value and in magnitude, essentially because the base value is positive and the largest correction comes from higher excitation ranks which is strongly positive.
The uncertainty on this final result is also determined from the presented series of calculations. An individual uncertainty for an effect is determined as the difference between the most accurate and the second most accurate model for that effect. Then, the uncertainties are $5$% for the atomic basis set, $70$% for higher excitation ranks, and $10$% for inner-shell correlation effects. From this, a conservative estimate of ±$\pm 80$% is derived for the final value, since the considered corrections partially act with opposing signs.
IV Conclusions
The presented variational relativistic configuration-interaction calculations of the hyperfine Stark coefficient are shown to be accurate in comparison with reliable literature results on the order of $<7$% for atomic 87Rb and 133Cs. The corresponding coefficient for 169Tm is significantly more difficult to calculate and the presented result is given with a consequently rather large uncertainty margin. The difficulties in obtaining accurate electron correlation effects for relevant properties of 169Tm have also been encountered in the calculation of its ground-state electric quadrupole moment Fleig_PhysRevA.107.032816 . Nevertheless, the hyperfine Stark coefficient for the ground-level hyperfine transition is smaller than $2.3\times 10^{-13}$ [Hz/((V/m)2)] – even in the $M_{J}=\frac{1}{2}$ levels where its value is greatest – which is about three orders of magnitude smaller than same coefficient for the hyperfine transitions in 87Rb and 133Cs. The present final result for the differential polarizability of the proposed clock transition in Ref. NatComm_Tm_2019 differs from the one measured and calculated in that paper by about a factor of 2. The origin of this discrepancy remains unclear and several sources of uncertainty in both results have to be addressed.
The presented method for the calculation of hyperfine Stark effects is well suited for future applications on other atomic systems, in particular those with clock transitions involving hole quantum states and/or several unpaired electrons.