### v2

# Variational Calculation of the Hyperfine Stark Effect in Atomic 87Rb, 133Cs, and 169Tm

Abstract
An electronically variational approach to the calculation of atomic hyperfine structure transition energies under the influence of static external electric fields is presented. The method avoids the calculation of intermediate atomic states entirely and requires only the wavefunctions of the electronic states involved in the respective hyperfine levels. These wavefunctions are obtained through relativistic general-excitation-rank configuration interaction theory. A variant of the method also enables for calculations on atoms with the most complicated of shell structures.
Applications to 87Rb, 133Cs and a specific clock transition in 169Tm are presented. The final results $k_{\text{Rb}}=-1.234\pm 0.0223$ $10^{-10}$ Hz/((V/m)2)] and $k_{\text{Cs}}=-2.347\pm 0.084$ $10^{-10}$ Hz/((V/m)2)] obtained under inclusion of up to quintuple excitations in the atomic wavefunction expansion are compatible with previous calculations and, in the case of Cs, confirm that one of the earlier experimental measurements is not reliable. For 169Tm that is used in the development of atomic clocks the differential static scalar electric dipole polarizability between ground levels $J=\frac{7}{2}$ and $J=\frac{5}{2}$ is calculated to be $\Delta\alpha^{s}_{0}=-0.134\pm 0.11$ a.u. This result from a pure ab initio calculation confirms the result of $\Delta\alpha^{s}_{0}=-0.063^{+0.01}_{-0.005}$ a.u. obtained in Nat. Comm. 10 (2019) 1724 where a combination of measurement and theoretical modeling has been used.
I Introduction
Among the various effects ludlow_rmp_2015 contributing to the frequency uncertainty of an atomic clock the blackbody radiation (BBR) shift plays a major role Itano_Wineland_Blackbody1982 ; porsev_derevianko_PRA2006 . The BBR shift, a systematic environmental perturbation, can be approximately related to the coefficient of the static electric-field shift of the clock transition Angstmann_Dzuba_Flambaum_PRL2006 ; Kozlov_Dzuba_Flambaum_PRA2014 . The latter coefficient, in turn, describes the difference between the static electric dipole polarizabilities of the respective clock-transition states Safronova_BBR_2009 . Their calculation typically requires a summation over a complete (bound and scattering) set of intermediate atomic states which can be problematic and has in the past led to disagreements (Angstmann_Dzuba_Flambaum_PRA2006 and references therein).
If the clock transition involves atomic hyperfine states then the static electric-field shift of those hyperfine levels in the external electric field schwartz_SPT_1959 ; Fortson_Ramsey_1964 needs to be determined. The present approach takes an external electric field into account variationally – i.e., to infinite order expressed in the language of perturbation theory – in the optimization of the atomic wavefunction. In this way, the summation over intermediate states Angstmann_Dzuba_Flambaum_PRA2006 ; beloy_derevianko_PRL2006 ; Safronova_UI_MS_2009 is avoided and only the field-dependent target-state electronic wavefunction of the atom has to be determined. The static Stark shift of a hyperfine clock transition then results from a simple expectation value of this wavefunction over the magnetic hyperfine operator.
Difficulties in addressing atomic quantum states with several open electronic shells (i.e., containing several unpaired electrons) and/or atomic states with a few holes in otherwise filled shells have been reported Safronova_BBR_2009 . The present approach and method has no such particular difficulties and the wavefunctions for more complicated quantum states of atoms can be obtained straightforwardly. This has recently been demonstrated in the calculation of electronic electric quadrupole moments of the thulium atom Fleig_PhysRevA.107.032816 which comprises an open $f$ shell in its electronic ground state.
Indeed, hyperfine-level transitions in the thulium atom have recently been used as clock transitions in the design of an optical atomic clock with unusually low sensitivity to the BBR NatComm_Tm_2019 ; PRA_Tm_2016 ; PRA_Tm_2020 . The relevant BBR frequency shift has been experimentally estimated NatComm_Tm_2019 to be a few orders of magnitude smaller than the corresponding shift in clock transitions of other neutral atoms. The uncertainty of this measurement has been given as around 5050% AIP_Tm_2020 for the electric component of the BBR. An extensive theoretical study of the thulium atom with relevance to its use as an atomic clock has been presented very recently fritzsche_Tm_2024 , but the BBR or hyperfine Stark shifts have not been addressed in that work.
The paper is organized as follows. In section II the theoretical approach is laid out in detail and a comparison with the common perturbative approaches to the calculation of hyperfine Stark shifts is drawn. In the following section III the hyperfine Stark coefficients for three atoms of relevance to atomic-clock research are presented using the present methods. In addition, the differential static dipole polarizability relevant to the clock transition in 169Tm is calculated using a finite-field approach that has earlier been applied to the calculation of spin-orbit resolved electric dipole polarizabilities in atomic states fleig_halo ; fleig_pol13 . In the final section IV conclusions from the presented work are drawn.
II Theory
The present approach differs substantially from previous approaches Angstmann_Dzuba_Flambaum_PRA2006 ; Safronova_BBR_2009 ; beloy_derevianko_PRL2006 to calculate electric polarizabilities of atoms including the hyperfine interaction. Before presenting the formal details it is, therefore, in place to lay out the big picture of the present method.
In a hypothetical atom without relativistic effects – in particular without the spin-orbit interaction – polarizabilities can be defined for orbital angular momentum ($L$) microstates and denoted $\alpha_{L,M_{L}}$. If now very weak spin-orbit interaction is taken into account then the total angular momentum $J$ becomes an exact quantum number. Under the assumption, however, that the magnetic coupling is very weak, $M_{L}$ remains an approximately valid quantum number and the corresponding polarizabilities can be labelled $\alpha_{J,L,(M_{L})}$. For clarity of notation, the approximate quantum number is set in parentheses.
By analogy, polarizabilities for atoms without magnetic hyperfine interaction but including the electronic spin-orbit interaction can be defined exactly as $\alpha_{J,M_{J}}$. Including now a very weak magnetic hyperfine interaction leaves $M_{J}$ as an approximately valid quantum number and the polarizabilities are correspondingly labelled $\alpha_{F,J,(M_{J})}$ where $F\in\{J+I,\ldots,|J-I|\}$ and $I$ is the nuclear spin quantum number. This is the picture adopted in the present approach.
In terms of methodology, the first step consists in variationally determining the wavefunctions of the atomic electronic states in question including relativistic effects, electron correlation effects, and fully relaxing the wavefunction with respect to the externally applied finite uniform electric field. In the second step, the obtained wavefunctions are used as zeroth-order wavefunctions in a perturbative first-order evaluation of the hyperfine interaction in the relevant atomic states.
II.1 Static Hyperfine Stark Shift
The following formulation is energy-based and in its present form only applicable to atomic $S$ (orbital angular momentum free) states where non-scalar static electric polarizability is zero. The more general case will be presented in forthcoming work. Thus, the present formulation is applied to Rubidium and Cesium electronic ground states. For Thulium, a different approach is chosen as to be discussed below.
We suppose a hyperfine doublet of states denoted with total angular-momentum quantum numbers $F_{u}$ (upper) and $F_{l}$ (lower) is subjected to a static external electric field $E_{\text{ext}}$. Treating the hyperfine interaction to first order in perturbation theory, the associated field-dependent transition energy can be written as
$\Delta\varepsilon(E_{\text{ext}})=\varepsilon_{J_{u},({M_{J}}_{u})}(E_{\text{ext}})+\varepsilon_{F_{u}}(E_{\text{ext}})-\big{(}\varepsilon_{J_{l},({M_{J}}_{l})}(E_{\text{ext}})+\varepsilon_{F_{l}}(E_{\text{ext}})\big{)}.$ (1)
where $J$ denotes total electronic angular momentum associated with the respective hyperfine level. $\varepsilon_{F}$ is thus the hyperfine energy relative to the respective electronic reference energy $\varepsilon_{J,M_{J}}$.
This transition energy can be related to a transition frequency which is also field dependent:
$\nu(E_{\text{ext}})=\frac{\Delta\varepsilon(E_{\text{ext}})}{h}$ (2)
where $h$ is Planck’s constant. The shift of this transition frequency due to $E_{\text{ext}}$ is in the present evaluated by using a finite but small $E$ field and a zero-field calculation. Then
$\delta\nu(E_{\text{ext}})=\nu(E_{\text{ext}})-\nu(E_{\text{ext}}=0).$ (3)
Following the definition in Ref. Safronova_BBR_2009 with according modifications and taking the electronic degrees of freedom into account the Stark coefficient $k$ as a function of this frequency shift is expressed as
$k=\frac{\delta\nu(E_{\text{ext}})}{E_{\text{ext}}^{2}}$ $=$ $\frac{1}{hE_{\text{ext}}^{2}}\left[\big{(}\varepsilon_{J_{u},({M_{J}}_{u})}(E_{\text{ext}})+\varepsilon_{F_{u}}(E_{\text{ext}})-\varepsilon_{J_{l},({M_{J}}_{l})}(E_{\text{ext}})-\varepsilon_{F_{l}}(E_{\text{ext}})\big{)}\right$. (5)
$\hskip 42.67912pt\left.-\big{(}\varepsilon_{J_{u},({M_{J}}_{u})}(0)+\varepsilon_{F_{u}}(0)-\varepsilon_{J_{l},({M_{J}}_{l})}(0)-\varepsilon_{F_{l}}(0)\big{)}\right]$
$=$ $\frac{1}{hE_{\text{ext}}^{2}}\left[\big{(}\varepsilon_{J_{u},({M_{J}}_{u})}(E_{\text{ext}})-\varepsilon_{J_{l},({M_{J}}_{l})}(E_{\text{ext}})\big{)}+\big{(}\varepsilon_{F_{u}}(E_{\text{ext}})-\varepsilon_{F_{l}}(E_{\text{ext}})\big{)}\right$.
$\hskip 42.67912pt\left.-\big{(}\varepsilon_{J_{u},({M_{J}}_{u})}(0)-\varepsilon_{J_{l},({M_{J}}_{l})}(0)\big{)}-\big{(}\varepsilon_{F_{u}}(0)-\varepsilon_{F_{l}}(0)\big{)}\right]$
and can, therefore, be calculated if the field-dependent level energies are known.
At this point a distinction between two cases can be made:
  1. 1.
    The two respective hyperfine levels belong to different electronic states. Then the Stark coefficient is stongly dominated by the electric polarizability difference between the different electronic wavefunctions and it can be written as
    $k\approx\frac{1}{hE_{\text{ext}}^{2}}\left[\big{(}\varepsilon_{J_{u},({M_{J}}_{u})}(E_{\text{ext}})-\varepsilon_{J_{l},({M_{J}}_{l})}(E_{\text{ext}})\big{)}-\big{(}\varepsilon_{J_{u},({M_{J}}_{u})}(0)-\varepsilon_{J_{l},({M_{J}}_{l})}(0)\big{)}\right]$ (6)
  2. 2.
    The two respective hyperfine levels belong to the same electronic state. In this case all of the electronic energies in the expression (5) cancel pairwise and the Stark coefficient is
    $k=\frac{1}{hE_{\text{ext}}^{2}}\left[\big{(}\varepsilon_{F_{u}}(E_{\text{ext}})-\varepsilon_{F_{l}}(E_{\text{ext}})\big{)}-\big{(}\varepsilon_{F_{u}}(0)-\varepsilon_{F_{l}}(0)\big{)}\right]$ (7)
II.2 Field-Dependent Hyperfine Level Energy
We will now be concerned with the above second case. A field-dependent hyperfine energy is related to the magnetic hyperfine constant $A$ as (see bethe_hypfin , p. 110)
$\varepsilon_{F}(E_{\text{ext}})\approx\frac{1}{2}\,\left[F(F+1)-I(I+1)-J(J+1)\right]\,A(E_{\text{ext}})$ (8)
where the quantum numbers $I$ and $J$ refer to nuclear spin and to total electronic angular momentum, respectively, of the state in question. The relationship Eq. (8) is not exact because the presence of the external field lifts the full rotational symmetry of the atom and thus also the strict validity of total angular momentum as a good quantum number. However, since the applied fields are very small (see below) the relation is still approximately correct.
Thus, if the field-dependent hyperfine constant $A(E_{\text{ext}})$ can be determined then also the Stark coefficient $k$. Following the implementation in Refs. Fleig2014 ; Fleig_Skripnikov2020 $A(E_{\text{ext}})$ is calculated as an expectation value
$A(E_{\text{ext}})=\left<\hat{H}_{\text{HF}}\right>_{\psi(E_{\text{ext}})}$ (9)
with the one-body hyperfine Hamiltonian
$\hat{H}_{\text{HF}}=-\frac{\mu[\mu_{N}]}{2cIm_{p}M_{J}}\,\sum\limits_{i=1}^{n}\,\left(\frac{\boldsymbol{\alpha}_{i}\times{\bf{r}}_{i}}{r_{i}^{3}}\right)_{z}$ (10)
over a field-dependent electronic wavefunction $\psi(E_{\text{ext}})$ where $n$ is the number of electrons, $\boldsymbol{\alpha}$ is a Dirac matrix, $\mu$ is the nuclear magnetic moment [in nuclear magnetons], $\frac{1}{2cm_{p}}$ is the nuclear magneton in a.u. , $m_{p}$ is the proton rest mass, and $\bf{r}$} is the electron position operator.
$\psi(E_{\text{ext}})$ is obtained by solving
$\hat{H}({{E}_{\text{ext}}})\big{|}\psi(E_{\text{ext}})\big{>}=\varepsilon({{E}_{\text{ext}}})\big{|}\psi(E_{\text{ext}})\big{>}$ (11)
with $\varepsilon({{E}_{\text{ext}}})$ the field-dependent energy eigenvalue and $\hat{H}({{E}_{\text{ext}}})$ the Dirac-Coulomb (DC) Hamiltonian including the interaction term with the external field:
$\hat{H}({{E}_{\text{ext}}})$ $:=$ $\hat{H}^{\text{Dirac-Coulomb}}+\hat{H}^{\text{Int-Dipole}}$ (12)
$=$ $\sum\limits^{n}_{j}\,\left[c\,\boldsymbol{\alpha}_{j}\cdot{\bf{p}}_{j}+\beta_{j}c^{2}-\frac{Z}{r_{jK}}{1\!\!1}_{4}\right]+\sum\limits^{n}_{k>j}\,\frac{1}{r_{jk}}{1\!\!1}_{4}+\sum\limits_{j}\,{\bf{r}}_{j}\cdot{\bf{E}_{\text{ext}}}\,{1\!\!1}_{4}$
${\bf{E}_{\text{ext}}}=E_{z}{\bf{e}}_{z}$ is uniform in space, the indices $j,k$ run over $n$ electrons, $Z$ is the proton number with the nucleus $K$ placed at the origin, and $\boldsymbol{\alpha},\beta$ are standard Dirac matrices. $E_{z}$ is not treated as a perturbation but included a priori in the variational optimization DIRAC_JCP of the wavefunction, $\psi(E_{z})$.
Technically, $\psi(E_{z})$ is a configuration interaction (CI) vector knecht_luciparII built from Slater determinants over field-dependent 4-spinors. This vector is represented as
[…]
where ${\cal{F}}^{t}(M,n)$ is the symmetry-restricted sector of Fock space with $n$ electrons in $M$ four-spinors, ${\cal{S}}=a^{\dagger}_{i}a^{\dagger}_{j}a^{\dagger}_{k}\ldots$ is a string of spinor creation operators, ${\overline{\cal{T}}}=a^{\dagger}_{\overline{l}}a^{\dagger}_{\overline{m}}a^{\dagger}_{\overline{n}}\ldots$ is a string of creation operators of time-reversal transformed spinors. The determinant expansion coefficients $c_{(\alpha,J,M_{J}),K}$ are generally obtained as described in refs. fleig_gasci ; fleig_gasci2 by diagonalizing the Dirac-Coulomb Hamiltonian.
II.3 Comparison with Perturbation Theory
In order to compare the present method with the perturbative approaches typically used in the literature, it is assumed that a zeroth-order CI problem for the target state $\psi_{0}$ is solved where $\hat{V}=\hat{H}^{\text{Int-Dipole}}=\sum\limits_{j}\,{\bf{r}}_{j}\cdot{\bf{E}_{\text{ext}}}\,{1\!\!1}_{4}$ is not included in the Hamiltonian $\hat{H}^{(0)}$:
$\hat{H}^{(0)}\big{|}\psi_{0}^{(0)}\big{>}=\varepsilon_{0}\big{|}\psi_{0}^{(0)}\big{>}$ (14)
Expanding the wavefunction $\left|\psi_{0}\right$> for the target state into a perturbation series where $\hat{V}$ is the perturbation and retaining only non-zero terms yields
$\left|\psi_{0}\right$> $=$ $\left|\psi_{0}^{(0)}\right>+\sum\limits_{k\neq 0}\,\frac{V_{k0}}{\varepsilon_{0}-\varepsilon_{k}}\left|\psi_{k}^{(0)}\right>-\frac{1}{2}\sum\limits_{k\neq 0}\,\frac{|V_{k0}|^{2}}{(\varepsilon_{0}-\varepsilon_{k})^{2}}\left|\psi_{0}^{(0)}\right$> (15)
$+\sum\limits_{k,l\neq 0}\,\frac{V_{kl}V_{l0}}{(\varepsilon_{0}-\varepsilon_{k})(\varepsilon_{0}-\varepsilon_{l})}\left|\psi_{k}^{(0)}\right>+{\cal{O}}(\hat{V}^{3})$
where $V_{kl}=\left<\psi_{k}^{(0)}|\hat{V}|\psi_{l}^{(0)}\right$> and $\varepsilon_{m}$ is an unperturbed energy. If terms ${\cal{O}}(\hat{V}^{3})$ are omitted from Eq. (15) and the resulting truncated expansion is used to evaluate the expectation value $\left<\psi_{0}\right|\hat{H}_{\text{HF}}\left|\psi_{0}\right$> then the third-order expressions from Refs. Angstmann_Dzuba_Flambaum_PRA2006 and schwartz_SPT_1959 arise. In the present work, however, ${\cal{O}}(\hat{V}^{3})$ is implicitly included in Eq. (11) which leads to the presence of all orders of the external electric field in Eq. (9) and, therefore, also in Eq. (5) defining the hyperfine Stark coefficient.
II.4 Scalar and Tensor Static Polarizabilites
An electric dipole polarizability for a state labelled $J,M_{J}$ can be given angel_sandars_pol1 ; Manakov_Ovsyannikov_1979 ; Chen_Raithel_2015 in terms of scalar ($\alpha_{0}$) and tensor ($\alpha_{2}$) static polarizabilites as follows:
$\alpha_{J,M_{J}}=\alpha_{0}(J)+\alpha_{2}(J)\frac{3M_{J}^{2}-J(J+1)}{J(2J-1)}$ (16)
where $\alpha_{0}$ and $\alpha_{2}$ are functions of $J$ only. If $\alpha_{J,M_{J}}$ and $\alpha_{J,M_{J}^{\prime}}$ with $M_{J}\neq M_{J}^{\prime}$ are known then the resulting system of Eqs. (16) can be inverted and $\alpha_{0}$ and $\alpha_{2}$ become calculable functions of $\alpha_{J,M_{J}}$ and $\alpha_{J,M_{J}^{\prime}}$. This fact will be exploited in order to compare with polarizability results given in terms of scalar and tensor polarizabilities in cases where $J>\frac{1}{2}$.
This is the case for the Thulium atom ground levels. In the calculations presented below $\alpha_{J,M_{J}}$ are calculated using the finite-field method (see subsection III.1.4). $\alpha_{0}(J)$ and $\alpha_{2}(J)$ are then obtained by inverting Eqs. (16). Finally, the differential static scalar polarizability in this case results from
$\Delta\alpha_{0}=\alpha_{0}(J_{u})-\alpha_{0}(J_{l})$ (17)
and the hyperfine Stark shift parameter can be obtained from this as
$k=-\frac{1}{2}\,\Delta\alpha_{0}$ (18)
III Applications and results
III.1 Technical Details
III.1.1 External Electric Field
The external electric field $E_{z}$ has to be chosen small enough to assure that $k$ as defined in Eq. (5) does not vary with $E_{z}$, but also large enough to ensure that the field-dependent change in $\psi(E_{z})$ is sufficiently greater than the convergence threshold chosen for the wavefunction. Typically, this is achieved for $E_{z}=10^{-3}$ a.u.  which is the value used in the present direct calculations. In addition, a region around $E_{z}=10^{-3}$ a.u.  has been explored which allows for an estimate of the uncertainty in $k$ due to deviations from quadratic dependency of the field-dependent total energy.
III.1.2 Atomic Basis Sets
For all three atoms Gaussian basis sets are used. The Rb atom is described by the uncontracted Dyall vTZ set with $28s,20p,12d,1f$ primitive functions and the aug-cc-pwCVQZ-X2C set from the EMSL library EMSL-basis2019 that has been singly densified in the $s$ and $p$ spaces and augmented by one dense and one diffuse function as described in Ref. Hubert_Fleig_2022 . The full uncontracted QZ+ set then comprises 77s,59p,21d,5f,3g77s,59p,21d,5f,3g functions. The Cs atom is described by the uncontracted Dyall vTZ set with $31s,24p,15d,1f$ primitive functions and the vQZ basis dyall_s-basis with $5s5p6s6p$ correlating exponents added. Like for Rb the $s$ and $p$ spaces have been densified once and augmented. The final uncontracted QZ+ set comprises $75s,61p,19d,3f,1g$ functions.
For the thulium atom two different basis sets dyall_4f are employed: 1) The uncontracted Dyall vTZ set including all $[4f/6s/5s,p,d]$ correlating and $4f$ dipole-polarizing functions adding up to $[30s,24p,18d,13f,4g,2h]$ functions (in the following denoted as TZ). 2) The uncontracted Dyall cvQZ basis including all $[4f/6s/5d]$ correlating and $4f$ dipole-polarizing functions comprised by a total of $[35s,30p,19d,16f,6g,4h,2i]$ functions (QZ).
The densification and augmentation in the $s$ and $p$ spaces in some cases serves to test a highly accurate basis set for describing the $s$-$p$ mixing which is predominantly responsible for the energy shifts when an external electric field is applied to hyperfine states.
III.1.3 Atomic Wavefunctions
All atoms considered here have an odd number of electrons. In addition, the interaction with $E_{z}$ breaks the full rotational symmetry of the atom which means that only $M_{J}$ remains as an exact electronic quantum number (for $\psi(E_{\text{ext}})$ in Eq. (11)). Atomic wavefunctions are calculated in the $\left|M_{J}\right|=\Omega$ irreducible representations of the $C_{\infty v}^{*}$ double point group. The target atom is placed at the origin of the reference frame and a ghost atom with neither electronic basis functions nor nuclear charge is placed at a finite distance along the axis of the external $E$ field in order to allow for the inclusion of the external-field Hamiltonian in linear-symmetry calculations. The ghost atom introduces no physical interaction.
For solving Eq. (11) the KRCI module knecht_luciparII of the DIRAC program package DIRAC_JCP is used. In a first step the Dirac-Coulomb-Hartree-Fock equations are solved where the Hamiltonian in Eq. (12) is employed. This model will be abbreviated as DCHF. The atomic spinors are optimized by diagonalizing a Fock operator where a fractional occupation of $f=\frac{m}{n}$ per spinor in the defined valence shells is used. Here $m$ is the number of electrons and $n$ is the number of spinors the respective shell comprises. These spinors are thus obtained for the electric potential of the neutral atom.
Acronyms are used for brevity in defining atomic correlated wavefunctions from the second step. As an example, SDT9_10au stands for Single, Double, Triple replacements relative to the DCHF reference state where 9 electrons occupying the outermost shells in the DCHF reference state are taken into account in the correlation expansion and the complementary space of virtual spinors is truncated at 10 a.u. An acronym SD8_SDT9 means that up to two holes are allowed in the shells occupied by 8 electrons and up to triple excitations into the virtual spinors are allowed from the combined shells occupied by the 9 electrons in the reference state. In the case of an alkali atom this means that up to double excitations from the (n1)s(n1)p(n-1)s~(n-1)p shells and up to triple excitations from the combined (n1)s(n1)p(n-1)s~(n-1)p and $ns$ shells are included in the wavefunction expansion. Thus, the model SD8_SDT9 comprises a subset of the determinants of the model SDT9, where in the latter all triple excitations from the (n1)s(n1)p(n-1)s~(n-1)p and $ns$ shells are included.
III.1.4 Finite-Field (FF) Method
Static dipole polarizabilities $\alpha_{J}$ are calculated by fitting finite-field electronic energies for four field points with $E\in\{0,0.00025,0.0005,0.001\}$ a.u. to a polynomial and extracting the second derivative of the fitted function at zero field which is proportional to the static dipole polarizability, see also Refs. fleig_halo ; fleig_pol13 .
III.2 87Rb
III.2.1 Static Electric Dipole Polarizability $\alpha_{D}$
As a corroboration of the present method of calculating polarizabilites a comparison with literature results for scalar and tensor polarizabilities is drawn. Using Eq. (16) the $M_{J}$-dependent values of $\alpha_{D}$ are obtained from the equations
$\alpha_{D}\left({}^{2}P_{3/2,1/2}\right)$ $=$ $\alpha_{0}\left({}^{2}P_{3/2}\right)-\alpha_{2}\left({}^{2}P_{3/2}\right)$
$\alpha_{D}\left({}^{2}P_{3/2,3/2}\right)$ $=$ $\alpha_{0}\left({}^{2}P_{3/2}\right)+\alpha_{2}\left({}^{2}P_{3/2}\right)$ (19)
where $\alpha_{0}$ is the scalar polarizability and $\alpha_{2}$ is the tensor polarizability. This allows for calculating the $M_{J}$-dependent polarizabilities from the results in Ref. Arora_Sahoo_2012 and these are given in Table 1. The inverted equations (19) read
$\alpha_{0}\left({}^{2}P_{3/2}\right)$ $=$ $\frac{1}{2}\left[\alpha_{D}\left({}^{2}P_{3/2,3/2}\right)+\alpha_{D}\left({}^{2}P_{3/2,1/2}\right)\right]$
$\alpha_{2}\left({}^{2}P_{3/2}\right)$ $=$ $\frac{1}{2}\left[\alpha_{D}\left({}^{2}P_{3/2,3/2}\right)-\alpha_{D}\left({}^{2}P_{3/2,1/2}\right)\right]$ (20)
This, in turn, allows for calculating the scalar and tensor polarizabilities from the present $M_{J}$-dependent values, the results of which are also given in Table 1.
Table 1: $M_{J}$-dependent static electric dipole polarizabilities $\alpha_{D}$ [a.u.] calculated through the FF method for states $^{M}L_{J,M_{J}}$ where $M=2S+1$ is the spin multiplicity and scalar ($\alpha_{0}$) and tensor ($\alpha_{2}$) polarizabilities for the $^{2}P_{3/2}$ state; for the electronic ground state $^{2}S_{1/2,1/2}$, $\alpha_{D}\left({}^{2}S_{1/2,1/2}\right)=\alpha_{0}\left({}^{2}S_{1/2}\right)$.
$5s^{1}$ $5p^{1}$
Model $^{2}S_{1/2,1/2}$ $^{2}P_{1/2,1/2}$ $^{2}P_{3/2,1/2}$ $^{2}P_{3/2,3/2}$ $\alpha_{0}(^{2}P_{3/2})$ $\alpha_{2}(^{2}P_{3/2})$
DCHF $485.3$
QZ+/SDT9/10au $333.0$ $820.3$ $1047.9$ $719.3$ $883.6$ 164.3-164.3
Experiment (cited in Ref. Arora_Sahoo_2012 ) $318.79$ $810.6$ $1020$ $694$ $857.0$ 163-163
Other theory (Arora_Sahoo_2012 ) $318.3$ $810.5$ $1033.9$ $702.1$ $868.0$ 165.9-165.9
For the $^{2}P_{3/2}$ level the present results for $\alpha_{0}$ and $\alpha_{2}$ which are derived from $M_{J}$-dependent calculations as described above differ from the experimental values by only 131-3%. Further improvements could be made by modifying the employed electronic-structure model, but for the present case the obtained correspondence serves as a sufficient proof of principle for the applied method.
III.2.2 Hyperfine Stark Effect
For the electronic ground state corresponding to the valence configuration $4s^{1}$ the total electronic angular momentum quantum number is J=1/2{J=1/2}. The considered isotope has $I=3/2$ and the resulting hyperfine quantum levels are denoted as $F_{u}=2$ and $F_{l}=1$. The fractional occupation in the DCHF calculation is $f=1/2$.
Results for the hyperfine Stark coefficient are shown in Table 2 for various electronic-structure models and are compared with experimental and theoretical literature results. As a general effect, the hyperfine constant of a given atom in an electronic $S$ state diminishes when an external $E$ field is included. This observation is explained by the fact that the $E$ field partially shifts spin density from $s$ wave to $p$ wave character in the atomic ground state (through $s$-$p$ mixing in the polarized atom), thus reducing the hyperfine interaction.
The hyperfine Stark coefficient in mean-field approximation (DCHF, both basis sets) differs from the experimental result by more than 2020% and lowest-order electron correlation effects from the $4s,4p,5s$ Rb shells even increase this discrepancy (model SD9, both basis sets). Upon including combined triple excitations from the $4s,4p$ shells and the $5s$ valence shell (model TZ/SD8_SDT9) a strong correction to $k$ is obtained. The inclusion of full triple excitations (model TZ/SDT9), i.e., including those from the $4s,4p$ shells, yields another large change in $k$. Close agreement with the result obtained by Safronova et al. Safronova_BBR_2009 is achieved at this level of calculation. This is not surprising since the wavefunction model used in Ref. Safronova_BBR_2009 for obtaining the cited result is a linearized coupled cluster expansion including up to (perturbative) valence triple excitations which is similar to the present SDT9 model.
Table 2: Stark hyperfine coefficient $k$ for 87Rb with nuclear magnetic moment $\mu=2.75131\mu_{N}$ ISOLDE_BWeffect_1993
Model $k$ $10^{-10}$ Hz/((V/m)2)]
TZ/DCHF 1.522-1.522
TZ/S8_SD9/10au 1.138-1.138
TZ/SD9/10au 1.618-1.618
TZ/SD8_SDT9/10au 1.220-1.220
TZ/SDT9/10au 1.273-1.273
TZ/SDT8_SDTQ9/10au 1.159-1.159
TZ/SDTQ9/10au 1.228-1.228
TZ/SDTQ8_SDTQQ9/10au 1.158-1.158
TZ/SDTQQ9/10au 1.161-1.161
TZ/SD18_SDT19/10au 1.253-1.253
TZ/SD26_SDT27/10au 1.295-1.295
TZ/SD34_SDT35/10au 1.299-1.299
TZ/SD36_SDT37/10au 1.301-1.301
QZ+/DCHF 1.546-1.546
QZ+/SD9/10au 1.662-1.662
QZ+/SD9/30au 1.669-1.669
QZ+/SD8_SDT9/30au 1.202-1.202
QZ+/SDT9/10au 1.265-1.265
QZ+/SDT9/30au 1.271-1.271
Final $\boldsymbol{-1.234\pm 0.0223}$}
Exp. Mowat_Li_1972 1.23(3)-1.23(3)
Safronova et al. Safronova_BBR_2009 1.272-1.272111preliminary value
Angstmann et al. Angstmann_Dzuba_Flambaum_PRA2006 1.24(1)-1.24(1)
It becomes clear that higher excitation ranks in the wavefunction expansion have a much greater impact on the hyperfine Stark coefficient than improvements in the atomic basis set. It is thus attempted to systematically converge $k$ with respect to the former effects. The evolution of the results for these systematically improved models is displayed graphically in Fig. 1 for ease of comparison.
Figure 1: Stark hyperfine coefficient $k$ for the electronic ground state of the rubidium atom using various electronic-structure models (black: present QZ+, blue: present TZ) and compared with other theoretical results and experiment, including the experimental uncertainty; for three of the models the evolution of the result with the number of CI iterations is also shown.
Along the improving model series SD9 – SD8_SDT9 – SDT9 – SDT8_SDTQ9 – SDTQ9 $\ldots$ results strongly oscillate, even after having included combined quadruple excitations. However, the partial series of corresponding models SD8_SDT9 – SDT8_SDTQ9 – SDTQ8_SDTQQ9 does lead to a sufficiently converged result. Likewise, the partial series SD9 – SDT9 – SDTQ9 – SDTQQ9 also reaches a sufficiently converged value that is not far from the result with the model SDTQ8_SDTQQ9. It can be inferred that the result from the model SDTQQ9 that includes full quintuple excitations for the $9$ outermost electrons is near the Full CI result in the TZ basis set.
The final value is obtained as follows. The result from the model SDTQQ9 serves as the base value and corrections from correlations of inner-shell electrons and from the atomic basis set are added. In detail,
$k({\text{final}})$ $=$ $k({\text{TZ/SDTQQ9/10au}})$ (21)
$+k({\text{TZ/SD36\_SDT37/10au}})-k({\text{TZ/SD8\_SDT9/10au}})$
$+k({\text{QZ+/SDT9/10au}})-k({\text{TZ/SDT9/10au}})$
This result differs from the experimental central value by only about 0.30.3% and is the theoretical result that comes closest to the experimental value. The uncertainty for the present final value is obtained by addition of individual uncertainties for interelectron correlation effects, the basis-set approximation and the approximation due to the use of the Dirac-Coulomb Hamiltonian operator (which neglects the Breit interaction and QED effects) for obtaining the atomic wavefunctions. The former two uncertainties are obtained from the difference between the most elaborate and the second most elaborate models, respectively. The latter uncertainty is estimated to be 1%.
III.3 133Cs
This important isotope of cesium has received considerable attention in the past. For the electronic ground state corresponding to the valence configuration $6s^{1}$ the total electronic angular momentum quantum number is J=1/2{J=1/2}. The considered isotope has $I=7/2$ and the resulting hyperfine quantum levels are denoted as $F_{u}=4$ and $F_{l}=3$. The fractional occupation in the DCHF calculation is $f=1/2$.
Results for 133Cs are compiled in Table 3. As in the case of 87Rb the mean-field result for $k$ is too large on the absolute, compared with the cited reference values, by around 4040%. Again, the inclusion of lowest-order electron correlation effects from the outermost atomic shells even increases this deviation. It is here in addition and for the model QZ+/SD9 shown that the consideration of virtual spinors of very high energy (up to $1000$ a.u. ) in the wave-function expansion does not affect the results significantly. Also as for 87Rb combined triple excitations from the outermost shells yield a very important correction to $k$. Including full triple excitations (model QZ+/SDT9/30au) yields a result that deviates from the most accurate literature results by roughly +9+9%. Accounting for combined quadruple excitations quenches the deviation to 7-7% where again, as in the Rb atom, the correction slightly overshoots the exact value. This latter model comprises about $1.9$ billion ($10^{9}$) terms in the CI expansion when the QZ+ basis set is used.
Table 3: Stark hyperfine frequency shift $k$ for 133Cs ($5s_{1/2}$), $I=3.5$, $E_{z}=0.001$ a.u., hyperfine level quantum numbers are $F_{u}=4$, $F_{l}=3$.
Model $k$ $10^{-10}$ Hz/((V/m)2)]
TZ/DCHF 3.164-3.164
TZ/SD9/10au 3.272-3.272
TZ/SD8_SDT9/10au 2.319-2.319
TZ/SDT9/10au 2.433-2.433
TZ/SDT8_SDTQ9/10au 2.153-2.153
TZ/SDTQ9/10au 2.321-2.321
TZ/SDTQ8_SDTQQ9/10au 2.156-2.156
TZ/SDTQQ9/10au 2.161-2.161
TZ/SD19/10au 3.335-3.335
TZ/SDT19/10au 2.516-2.516
TZ/SDT27/10au 2.595-2.595
TZ/SD27/10au 3.395-3.395
TZ/SD37/10au 3.396-3.396
TZ/SD45/10au 3.380-3.380
QZ+/DCHF 3.220-3.220
QZ+/SD9/10au 3.375-3.375
QZ+/SD9/30au 3.379-3.379
QZ+/SD9/1000au 3.384-3.384
QZ+/SD8_SDT9/10au 2.308-2.308
QZ+/SD8_SDT9/30au 2.311-2.311
QZ+/SDT9/10au 2.468-2.468
QZ+/SDT9/30au 2.469-2.469
QZ+/SDT8_SDTQ9/10au 2.112-2.112
Final $\boldsymbol{-2.347\pm 0.084}$}
Exp.Simon_PRA1998 2.271(4)-2.271(4)
Exp.Godone_PRA2005 2.05(5)-2.05(5)
Safronova et al.beloy_derevianko_PRL2006 2.271(8)-2.271(8)
Angstmann et al.Angstmann_Dzuba_Flambaum_PRA2006 2.26(2)-2.26(2)
Even higher excitation ranks can with the present method only be treated in the smaller TZ basis set. There, even a calculation with full quadruple excitations does not yet deliver a converged result (model TZ/SDTQ9). A converged result, however, is obtained when at least combined quintuple excitations are included in the wavefunction expansion.
Fig. 2 displays a selection of results supporting the present discussion. Correlations among and with the electrons from inner shells down to and including the $3s$ shell (model SD45) lead to a small increase of $k$, on the absolute. A very similar correction is obtained when full triple excitations are included in these expansions (models SDT9 and SDT27). This suggests that there is a negligible effect from including even higher excitation ranks in order to obtain corrections from inner-shell electron correlations.
Figure 2: Stark hyperfine coefficient $k$ for the electronic ground state of the cesium atom using various electronic-structure models (black: present QZ, blue: present TZ) and compared with other theoretical results and experiment, including the experimental uncertainty; for two of the models the evolution of the result with the number of CI iterations is also shown.
The final present value of the hyperfine Stark shift for Cs is, therefore, obtained as follows. The base value is taken from the model TZ/SDTQ8_SDTQQ9. This result is corrected by inner-shell correlations from the $4s,4p,4d$ shells at the SDT level and for inner-shell correlations from the $3s,3p,3d$ shells at the SD level. A small basis-set correction is added from the SDT9 model as well as a correction from including virtual spinors up to 1000 a.u. (model QZ+/SD9). Mathematically, the corresponding evaluations are
$k({\text{final}})$ $=$ $k({\text{TZ/SDTQ8\_SDTQQ9/10au}})$ (22)
$+k({\text{TZ/SDT27/10au}})-k({\text{TZ/SDT9/10au}})$
$+k({\text{TZ/SD45/10au}})-k({\text{TZ/SD27/10au}})$
$+k({\text{QZ+/SD9/1000au}})-k({\text{QZ+/SD9/10au}})$
$+k({\text{QZ+/SDT9/10au}})-k({\text{TZ/SDT9/10au}})$
The uncertainty on this final result is estimated in the same way as has been done for the above Rb atom.
The present final result is compatible with the measurement by Simon et al. from 1998 Simon_PRA1998 and incompatible with the measurement by Godone et al. from 2005 Godone_PRA2005 . The present result is also compatible with the theoretical results by Safronova et al. beloy_derevianko_PRL2006 and by Angstmann et al. Angstmann_Dzuba_Flambaum_PRA2006 . The result by Godone et al. is in conflict with all other experimental and theoretical results. The present result supports previous theoretical values as well as the measurement by Simon et al.
III.4 169Tm
With confidence in the method established in the aforegoing sections it is now applied to an atom where high-level theoretical reference results are not available and which has a more complex electronic structure. In this case the fractional occupation in the DCHF calculation is $f=15/16$ where $m=15$ represents the thirteen $4f$ electrons plus the two $6s$ electrons. This averaging was required to assure proper convergence of the DCHF wavefunction.
III.4.1 $^{2}F_{7/2}(F=4)$$^{2}F_{5/2}(F=3)$ clock transition
The clock transition discussed in Refs. NatComm_Tm_2019 ; PRA_Tm_2016 ; PRA_Tm_2020 comprises a hyperfine component ($F_{l}=4$) of the ground electronic state $^{2}F_{7/2}$ and a hyperfine component ($F_{u}=3$) of the first electronically excited state $^{2}F_{5/2}$. In this case the Stark coefficient is dominated by the polarizability difference between the two respective electronic states and Eq. (17) applies.
Individual static polarizabilities $\alpha_{J}$ are calculated by using the FF method and are given in Table 4. Using Eq. (16) the $M_{J}$-dependent values of $\alpha_{D}$ are in the case of thulium states with $J=7/2$ obtained from the inverted equations
$\alpha_{0}\left({}^{2}F_{7/2}\right)$ $=$ $\frac{7}{6}\alpha_{D}\left({}^{2}F_{7/2,5/2}\right)-\frac{1}{6}\alpha_{D}\left({}^{2}F_{7/2,7/2}\right)$ (23)
$\alpha_{2}\left({}^{2}F_{7/2}\right)$ $=$ $\frac{7}{6}\left[\alpha_{D}\left({}^{2}F_{7/2,7/2}\right)-\alpha_{D}\left({}^{2}F_{7/2,5/2}\right)\right]$ (24)
$\alpha_{0}\left({}^{2}F_{7/2}\right)$ $=$ $\frac{5}{2}\alpha_{D}\left({}^{2}F_{7/2,3/2}\right)-\frac{3}{2}\alpha_{D}\left({}^{2}F_{7/2,1/2}\right)$ (25)
$\alpha_{2}\left({}^{2}F_{7/2}\right)$ $=$ $\frac{7}{2}\left[\alpha_{D}\left({}^{2}F_{7/2,3/2}\right)-\alpha_{D}\left({}^{2}F_{7/2,1/2}\right)\right]$ (26)
results for which are also given in Table 4.
Table 4: $M_{J}$-dependent static electric dipole polarizabilities $\alpha_{D}$ [a.u.] calculated through the FF method for states $^{M}L_{J,M_{J}}$ where $M=2S+1$ is the spin multiplicity and scalar ($\alpha_{0}$) and tensor ($\alpha_{2}$) polarizabilities for the $^{2}F_{7/2}$ multiplet
$\alpha_{D}$
Model $^{2}F_{7/2,1/2}$ $^{2}F_{7/2,3/2}$ $^{2}F_{7/2,5/2}$ $^{2}F_{7/2,7/2}$ $\alpha_{0}(^{2}F_{7/2})$ $\alpha_{2}(^{2}F_{7/2})$ calc. from $\{M_{J}\}$
TZ/DCHF $186.4554$ $185.7002$ $184.5674$ 2.643-2.643 $1/2,3/2$
TZ/SD15/6au $163.9967$ $163.2578$ $162.1495$ 2.586-2.586 $1/2,3/2$
TZ/SD13_SD15/6au $163.9897$ $163.2515$ $162.1442$ 2.584-2.584 $1/2,3/2$
TZ/SD13_SDT15/6au $161.9754$ $161.3215$ $160.3407$ 2.289-2.289 $1/2,3/2$
TZ/SD13_SDTsppdQ15/6au $167.2591$ $166.5212$ $165.4144$ 2.583-2.583 $1/2,3/2$
QZ/DCHF $186.6210$ $185.8505$ $184.6948$ 2.697-2.697 $1/2,3/2$
QZ/SD15/10au $161.2430$ $160.4891$ $158.9842$ $156.7269$ $159.3583$ 2.639-2.639 $1/2,3/2$
$159.3604$ 2.634-2.634 $5/2,7/2$
QZ/SDT15/10au $160.8884$ $160.2111$ $159.1952$ 2.371-2.371 $1/2,3/2$
QZ/SD23/10au $151.2078$ $150.4219$ $149.2431$ 2.751-2.751 $1/2,3/2$
QZ/SD33/10au $151.8781$ $151.0859$ $149.5087$ $147.1326$ $149.8976$ 2.773-2.773 $1/2,3/2$
$149.9047$ 2.772-2.772 $5/2,7/2$
Experiment $130\pm 16$ Ma_Tl_pol
Recommended $144\pm 15$ schwerdt_nagle_pol_atoms_2018
Both the scalar and tensor polarizabilites are numerically invariant to the choice of $M_{J}$ components used for their calculation up to two digits after the decimal point. This is shown for one case using the model QZ/SD33/10au. All scalar and tensor polarizabilites are therefore calculated from $\alpha_{7/2,3/2}$ and $\alpha_{7/2,1/2}$.
For $J=5/2$ states of thulium the inverted equations read
$\alpha_{0}\left({}^{2}F_{5/2}\right)$ $=$ $\frac{4}{3}\alpha_{D}\left({}^{2}F_{5/2,3/2}\right)-\frac{1}{3}\alpha_{D}\left({}^{2}F_{5/2,1/2}\right)$ (27)
$\alpha_{2}\left({}^{2}F_{5/2}\right)$ $=$ $\frac{5}{3}\left[\alpha_{D}\left({}^{2}F_{5/2,3/2}\right)-\alpha_{D}\left({}^{2}F_{5/2,1/2}\right)\right]$ (28)
$\alpha_{0}\left({}^{2}F_{5/2}\right)$ $=$ $\frac{1}{6}\alpha_{D}\left({}^{2}F_{5/2,5/2}\right)+\frac{5}{6}\alpha_{D}\left({}^{2}F_{5/2,3/2}\right)$ (29)
$\alpha_{2}\left({}^{2}F_{5/2}\right)$ $=$ $\frac{5}{6}\left[\alpha_{D}\left({}^{2}F_{5/2,5/2}\right)-\alpha_{D}\left({}^{2}F_{5/2,3/2}\right)\right]$ (30)
This, in turn, allows for calculating the scalar and tensor polarizabilities from the present $M_{J}$-dependent values, the results of which are also given in Table 5.
Table 5: $M_{J}$-dependent static electric dipole polarizabilities $\alpha_{D}$ [a.u.] calculated through the FF method for states $^{M}L_{J,M_{J}}$ where $M=2S+1$ is the spin multiplicity and scalar ($\alpha_{0}$) and tensor ($\alpha_{2}$) polarizabilities for the $^{2}F_{5/2}$ multiplet
$\alpha_{D}$
Model $^{2}F_{5/2,1/2}$ $^{2}F_{5/2,3/2}$ $^{2}F_{5/2,5/2}$ $\alpha_{0}(^{2}F_{5/2})$ $\alpha_{2}(^{2}F_{5/2})$ calc. from $\{M_{J}\}$
TZ/DCHF $186.3443$ $185.0430$ $184.6092$ 2.169-2.169 $1/2,3/2$
TZ/SD15/6au $163.9999$ $162.7315$ $162.3087$ 2.114-2.114 $1/2,3/2$
TZ/SD13_SD15/6au $163.9924$ $162.7252$ $162.3028$ 2.112-2.112 $1/2,3/2$
TZ/SD13_SDT15/6au $161.9244$ $160.8021$ $160.4280$ 1.871-1.871 $1/2,3/2$
TZ/SD13_SDTsppdQ15/6au $167.1381$ $165.8626$ $165.4374$ 2.126-2.126 $1/2,3/2$
QZ/DCHF $186.5082$ $185.1807$ $184.7382$ 2.213-2.213 $1/2,3/2$
QZ/SD15/10au $161.2017$ $159.9065$ $159.4748$ 2.159-2.159 $1/2,3/2$
QZ/SDT15/10au $160.7452$ $159.5866$ $159.2004$ 1.931-1.931 $1/2,3/2$
QZ/SD23/10au $151.1002$ $149.7506$ $149.3007$ 2.249-2.249 $1/2,3/2$
QZ/SD33/10au $151.7513$ $150.3923$ $147.6753$ $149.9393$ 2.265-2.265 $1/2,3/2$
QZ/SD33/10au $149.9395$ 2.264-2.264 $5/2,3/2$
In Ref. NatComm_Tm_2019 Eq. (5) is given
$\Delta\alpha_{\text{DC}}^{s}=\alpha^{s}_{5/2}-\alpha^{s}_{7/2}=-0.063(30)~{\text{a.u.}}$ (31)
from a combination of measurement and calculation. This being a negative quantity means that $\alpha^{s}_{7/2}>\alpha^{s}_{5/2}$.
The electronic ground state $^{2}F_{7/2}$ of the Tm atom can in the Hartree-Fock picture be represented by a $4f^{13}$ configuration written as $4f_{5/2}^{6}\,4f_{7/2}^{7}$ in terms of Hartree-Fock spinors (the $j=5/2$ level is energetically lower than the $j=7/2$ level). In turn, the first excited state $^{2}F_{5/2}$ can be represented as $4f_{5/2}^{5}\,4f_{7/2}^{8}$. According to Ref. desclaux_tables numerical Dirac-Hartree-Fock calculations show that the radial expectation values of the valence spinors are
$\left<\hat{r}\right>_{5/2}$ $=$ $0.763~{\text{a.u.}}$
$\left<\hat{r}\right>_{7/2}$ $=$ $0.780~{\text{a.u.}}$
respectively. This means that, qualitatively, the level with the greater $4f_{7/2}$ occupation is in a straightforward interpretation expected to be the level with the greater static dipole polarizability, since the $4f_{7/2}$ spinors are more diffuse than the $4f_{5/2}$ spinors. The level with the greater $4f_{7/2}$ occupation is the $^{2}F_{5/2}$ level ($8$ electrons occupying the $j=7/2$ spinors). Therefore, in Dirac-Hartree-Fock theory, the expected result is $\alpha^{s}_{5/2}>\alpha^{s}_{7/2}$ which contradicts the result given in Ref. NatComm_Tm_2019 and in Eq. 31 above.
Relativistic many-body calculations yield results for the differential scalar static dipole polarizability defined as $\Delta\alpha_{0}=\alpha_{0}(^{2}F_{5/2})-\alpha_{0}(^{2}F_{7/2})$ and compiled in Table 6.
Table 6: Differential static scalar electric dipole polarizabilities $\alpha_{0}$ [a.u.] for the thulium atom ground term calculated through the FF method from various CI models
Model $\alpha_{0}(^{2}F_{5/2})$ $\alpha_{0}(^{2}F_{7/2})$ $\Delta\alpha_{0}$
TZ/DCHF/0au $184.6092$ $184.5674$ $0.0418$
TZ/SD15/6au $162.3087$ $162.1495$ $0.1592$
TZ/SD13_SD15/6au $162.3028$ $162.1442$ $0.1586$
TZ/SD13_SDT15/6au $160.4280$ $160.3407$ $0.0873$
TZ/SD13_SDTsppdQ15/6au $165.4374$ $165.4144$ $0.0230$
QZ/DCHF/0au $184.7382$ $184.6948$ $0.0434$
QZ/SD15/10au $159.4748$ $159.3583$ $0.1165$
QZ/SDT15/10au $159.2004$ $159.1952$ $0.0052$
QZ/SD23/10au $149.3007$ $149.2431$ $0.0576$
QZ/SD33/10au $149.9393$ $149.8976$ $0.0417$
final $\boldsymbol{-0.134\pm 0.11}$}
Experiment NatComm_Tm_2019 0.063(30)-0.063(30)
It is interesting to note that at all individual levels of calculation the qualitative picture of DCHF theory is reproduced also when electron correlation effects are taken into account. However, when adding the effect of Triple excitations to the model with the greatest number of electrons subjected to the correlation treatment, SD33/10au, the mentioned interpretation at Hartree-Fock level of theory is reverted. Another sizable negative correction from a limited set of quadruple excitations on top of the triple excitations is determined by using the smaller TZ basis set where a calculation of this size becomes feasible. Going beyond this model is not possible with the current implementation due to computational limitations and calculation time.
The final result for the differential static scalar dipole polarizability for the Tm atom $^{2}F_{7/2}(F=4)$$^{2}F_{5/2}(F=3)$ clock transition is obtained by using the result from the model with the greatest number of electrons in the correlation treatment (SD33/10au) as a base value and adding to it corrections due to CI excitation ranks surpassing singles and doubles excitations. In formal terms this calculation reads as
$\Delta\alpha_{0}({\text{final}})$ $=$ $\Delta\alpha_{0}({\text{QZ/SD33/10au}})$
$+\Delta\alpha_{0}({\text{QZ/SDT15/10au}})-\Delta\alpha_{0}({\text{QZ/SD15/10au}})$
$+\Delta\alpha_{0}({\text{TZ/SD13\_SDTsppdQ15/6au}})-\Delta\alpha_{0}({\text{TZ/SD13\_SDT15/6au}})$
Physically speaking, an encompassing treatment of interelectron correlation effects leads to the astonishing result that the static scalar electric dipole polarizability is greater in the $^{2}F_{7/2}$ state than in the $^{2}F_{5/2}$ state of thulium. The simple single-determinant picture for electric polarizability based on Hartree-Fock spinors, therefore, breaks down. The present result confirms qualitatively in this regard the result from Ref. NatComm_Tm_2019 that has been obtained through a combination of experimental measurement and theory. However, due to the sizable corrections found with even the most extensive CI models the uncertainty on the present result must remain rather large. The uncertainty based on the limited treatment of CI excitation ranks, atomic basis sets, correlation effects from inner atomic shells and approximations in the employed atomic Hamiltonian does not compromise the qualitative conclusions.
IV Conclusions
A variational relativistic configuration-interaction approach to the calculation of the hyperfine Stark coefficient in atoms is presented. As a methodological conclusion, the present approach can be applied to electronic transitions of any type in any atom, given that the hyperfine Stark coefficient is calculated as a differential static polarizability.
For the 87Rb atom excitation ranks up to quintuples have been included in the wavefunction expansion and the obtained central value is the theoretical result so far closest to the experimental value. Calculations of similar sophistication for the 133Cs atom yield a final result that is compatible with other high-level theoretical calculations and the 1998 experimental value by Simon et al. Simon_PRA1998 . The present method is then applied to a 169Tm clock transition where so far pure ab-initio calculations have been lacking. The present calculations explain how the sign of the hyperfine Stark coefficient that has previously been measured NatComm_Tm_2019 comes about. The difficulties in obtaining accurate electron correlation effects for relevant properties of 169Tm have also been encountered in the calculation of its ground-state electric quadrupole moment Fleig_PhysRevA.107.032816 .